# A Visual-Inertial Motion Prior SLAM for Dynamic Environments

### v1

@@ Abstract
The Visual-Inertial Simultaneous Localization and Mapping (VI-SLAM) algorithms which are mostly based on static assumption are widely used in fields such as robotics, UAVs, VR, and autonomous driving. To overcome the localization risks caused by dynamic landmarks in most VI-SLAM systems, a robust visual-inertial motion prior SLAM system, named (IDY-VINS), is proposed in this paper which effectively handles dynamic landmarks using inertial motion prior for dynamic environments to varying degrees. Specifically, potential dynamic landmarks are preprocessed during the feature tracking phase by the probabilistic model of landmarks’ minimum projection errors which are obtained from inertial motion prior and epipolar constraint. Subsequently, a bundle adjustment (BA) residual is proposed considering the minimum projection error prior for dynamic candidate landmarks. This residual is integrated into a sliding window based nonlinear optimization process to estimate camera poses, IMU states and landmark positions while minimizing the impact of dynamic candidate landmarks that deviate from the motion prior. Finally, experimental results demonstrate that our proposed system outperforms state-of-the-art methods in terms of localization accuracy and time cost by robustly mitigating the influence of dynamic landmarks.
@@ I INTRODUCTION
In environments without satellite positioning, V-SLAM obtains information of the surrounding environment through visual sensors to solve its own motion. It has the advantages of low cost, high precision in trajectory estimation within small areas, no need for any environmental prior, and rich applicable scenarios. However, visual-only sensor is easily affected by external environmental and motion factors, such as dynamic landmarks, the lack of texture, lighting conditions, motion blur, and few common field of view (FOV) [1]. Therefore, an inertial measurement unit (IMU), which can provide relatively accurate pose estimation information without being affected by the environment in a short period of time, can be integrated to compensate for the deficiencies of visual-only sensor [2].
However, the assumption of static landmarks in visual-related SLAM still makes it has hidden localization risk in environments with dynamic landmarks [3]. The dynamic landmarks will affect the accuracy of visual positioning, and in extreme cases, they will lead to rapid deterioration of positioning.
@@
(2) The bundle adjustment (BA) residual for dynamic candidate landmarks: A dynamic visual residual is constructed which is used in nonlinear optimization process based on sliding window to reduce the impact of landmarks that deviate from the motion prior.
Compared with existing methods in the VIODE dataset [8] which contains dynamic objects within different ranges, superior robust and stable localization accuracy, and good real-time performance of our approach are demonstrated. In addition, the localization accuracy is compared in the EUROC dataset [9], and the result demonstrates stable and superior performance of our approach in static environments.
@@ II-C Methods based on inertial sensor
Cui et al. [22] detected and removed dynamic landmarks according to the epipolar constraint provided by the prior of IMU pre-integration, and conducted states estimation by utilizing more accurate static landmarks. The DynaVINS [23] proposed a robust bundle adjustment system based on the IMU pre-integration regularization factor and the weight momentum factor to reject dynamic landmarks. The method in this paper is also based on the motion prior propagated by the IMU which is used to label potential dynamic landmarks, and we propose a robust bundle adjustment residual for dynamic candidate landmarks.
@@ III-B Minimum Projection Error and Preprocessing
represents the precision of the epipolar constraint, is the coordinate in the normalized camera coordinate system while is the coordinate in the pixel coordinate system, is the coordinate measurement combined the true value and the residual as Fig.2 (b), is the internal parameter matrix of camera.
@@
is the minimum square projection error defined by us. is a factor that characterizes the credibility of the motion prior obtained through IMU propagation, which is defined to avoid the incorrect elimination of dynamic landmarks caused by large IMU propagation error (mainly considered as translation error) or the degradation of the epipolar constraint due to an overly small translation amount. is the variance of the position error of IMU pre-integration. Noted that for the pinhole camera model, we adopt the projection error on the pixel plane, while for the fisheye camera model, we use the projection unit length error on the normalized plane.
@@
(1) Elimination in preprocess: In the situation where both dynamic and static landmarks exist, the value of CV is always greater than . Sort all landmarks in descending order according to the value of , and start eliminating from the maximum value until the CV no longer decreases or the CV decreases to the target value . The elimination criterion and the recursive formulas for mean, variance and CV of remaining landmarks are as follows.
@@
(3) Mask for extracting landmarks: After elimination and marking, the area within around the eliminated dynamic landmark is considered the dynamic candidate area and no new landmarks will be extracted.
@@ IV ROBUST BUNDLE ADJUSTMENT
When a dynamic candidate landmark enters the sliding window based optimization process, it is continuously observed for times. The visual residual for these observations is as which considers the prior of minimum projection error which will also be updated during the iterative optimization process. From the probabilistic perspective, the two-dimensions’ projection residual of dynamic candidate landmarks satisfies a Gaussian distribution with a covariance matrix of minimum value as .
represents the collection of dynamic candidate landmarks in the sliding window, represents the observation to landmark , represents the camera pose state, IMU state and landmarks state that are to be optimized. and is the reduction factor of dynamic candidate landmarks’ visual residuals, and it is related to the prior of minimum projection error (more than one ) obtained through the epipolar constraint precision between the first observation and the observation in the sliding window. is huber kernel function [26]. Noted that .
Combined with the visual residuals of static landmarks, the measurement residuals of the IMU and the prior residual, the global residual to be minimized is (8).
is the collection of static landmarks in the sliding window, and is the covariance matrix of static landmark’s projection residual. represents the collection of inertial measurements while is the residual covariance matrix of the IMU pre-integration [24] between the and frames. represents prior information.
After optimization, 3D-2D projection from the first observation to other relevant observations is performed to calculate the reprojection errors of dynamic candidate landmarks. Remove them outside the sliding window and mark them as confirmed dynamic landmarks if the errors are still relatively large.
@@ V EXPERIMENTAL RESULTS
To evaluate our proposed algorithm, comparisons with the SOTA algorithms in the VIODE and EUROC datasets are carried out, such as VINS-FUSION [14], VINS-FUSION_ransac and Dyna-VINS [23] in the stereo mode.
@@ V-B Remove dynamic landmarks
After optimization based on sliding window using our visual residual for dynamic candidate landmarks and post-processing to them, confirmed dynamic landmarks will be removed from sliding window while others will be marked as static landmarks. The reduced dynamic visual residual result compared with original residual is displayed in Fig.5 and final labels of landmarks in Fig.3 is displayed in Fig.7.
V-C Qualitative Analyses
@@
This subsection provides a quantitative analysis of our method, evaluating the localization accuracy of camera poses and the time taken. Tables I and II present the comparison of our proposed method with other state-of-the-art SLAM methods, using the RMSE (Root Mean Square Error) of ATE (Absolute Trajectory Error) and average tracking and optimization time cost as the evaluation metric. To comprehensively validate the fitness for environments within different dynamic ranges and the effectiveness of the proposed method, we conducted tests on the VIODE dynamic dataset and the EUROC static dataset.
@@
Table I Comparison of Localization accuracy with state-of-the-art methods (RMSE of ATE in [m]). Noted that x represents failure case (diverged).
| Method | VIODE city | VIODE parking_lot | EUROC | ||||||||
| none | low | mid | high | none | low | mid | high | V1_03 | V2_03 | MH_03 | |
| VINS-Fusion | 0.158 | 0.169 | 0.193 | 0.211 | 0.103 | 0.126 | 0.149 | 0.166 | 0.216 | 0.352 | 0.290 |
| VINS-Fusion_ransac | 0.318 | 0.289 | 0.298 | 0.301 | 0.099 | 0.097 | x | 0.310 | 0.109 | 0.370 | 0.266 |
| Dyna-VINS | 0.157 | 0.140 | 0.149 | 0.122 | 0.115 | 0.099 | 0.126 | 0.124 | 0.175 | 0.479 | 0.349 |
| IDY-VINS | 0.148 | 0.145 | 0.135 | 0.126 | 0.092 | 0.089 | 0.111 | 0.088 | 0.143 | 0.229 | 0.276 |
@@
The effects of the proposed method on the localization accuracy is analyzed as shown in Table I. And Table II shows the analysis of our method on the tracking and optimization time cost. The comparison of the trajectory results in the directions for different methods and ground truth are shown in Fig.8.
@@
(1) The improvement of localization accuracy for different environments: According to Table I, our method has a better result than VINS-Fusion and Dyna-VINS while the RANSAC method has the problem of erroneously removing static inliers when facing large-scale dynamic scenes, like Fig.3(a), which leads to the decline and instability of the accuracy, although the RANSAC method can achieve good results in some cases. For outdoor sequences of VIODE from day_none to day_high, our method improves the localization accuracy by 6.3%, 14.2%, 30.1% and 40.3% compared to VINS-Fusion. And our method improves the localization accuracy by 10.7%, 29.4%, 25.5% and 47.0% compared to VINS-Fusion for indoor parking lot sequences of VIODE from parking_lot_none to parking_lot_high. The results of EUROC sequences in real world prove that our method still has a stable improvement effect on the localization accuracy in static environments. Although in some cases, the improvement effect is not as good as RANSAC method for EUROC dataset, but it is better than Dyna-VINS method. Moreover, the performance of Dyna-VINS method is related to complex parameters.
@@
| IDY-VINS | 20.403 | 20.759 | 20.628 | 20.908 | 38.375 | 37.375 | 37.251 | 35.564 | ||||
(2) Better time cost performance: According to Table II, our method does not cause too much additional tracking time cost compared to RANSAC method. However, it should be noted that the time consumption will surely be higher than VINS-Fusion and Dyna-VINS which do not conduct additional elimination operations during the feature tracking phase. On the other hand, the optimized results converge more quickly for our method than VINS-Fusion and VINS-Fusion_ransac methods because of the elimination of obvious dynamic landmarks during feature tracking and the classification of dynamic candidate landmarks and static landmarks during optimization process. And the more dynamic objects there are in the environment, the better improvement on the consumption of optimization time will be for our method, while the RANSAC method often requires a longer optimization time to achieve convergence with a lower localization accuracy because of incorrect elimination. Meanwhile, Dyna-VINS method requires much more time for optimization due to its complex processing. The above two aspects of tracking time cost and optimization time cost demonstrate that our method generally has a better real-time performance.
@@ VI CONCLUSIONS
In this paper, the robust visual-inertial motion prior SLAM system (IDY-VINS) has been proposed, which effectively handles dynamic landmarks using inertial motion prior for dynamic environments to varying degrees. It is a robust system that preprocesses potential dynamic landmarks and constructs bundle adjustment (BA) residual for dynamic candidate landmarks to reduce their impact. Experimental evidences demonstrate that our algorithm has a superior performance compared to other algorithms in both simulated and real environments, both dynamic and static environments.
In future work, we plan to extend the consideration of dynamic landmarks and candidate landmarks to the overall relocalization and loop closure detection processes of VI-SLAM.

### v2

@@ Abstract
The Visual-Inertial Simultaneous Localization and Mapping (VI-SLAM) algorithms which are mostly based on static assumption are widely used in fields such as robotics, UAVs, VR, and autonomous driving. To overcome the localization risks caused by dynamic landmarks in most VI-SLAM systems, a robust visual-inertial motion prior SLAM system, named (IDY-VINS), is proposed in this paper which effectively handles dynamic landmarks using inertial motion prior for dynamic environments to varying degrees. Specifically, potential dynamic landmarks are preprocessed during the feature tracking phase by the probabilistic model of landmarks’ minimum projection errors which are obtained from inertial motion prior and epipolar constraint. Subsequently, a robust and self-adaptive bundle adjustment residual is proposed considering the minimum projection error prior for dynamic candidate landmarks. This residual is integrated into a sliding window based nonlinear optimization process to estimate camera poses, IMU states and landmark positions while minimizing the impact of dynamic candidate landmarks that deviate from the motion prior. Finally, a clean point cloud map without ‘ghosting effect’ is obtained that contains only static landmarks. Experimental results demonstrate that our proposed system outperforms state-of-the-art methods in terms of localization accuracy and time cost by robustly mitigating the influence of dynamic landmarks.
@@ I INTRODUCTION
In environments without satellite positioning, V-SLAM obtains information of the surrounding environment through visual sensors to solve its own motion. It has the advantages of low cost, high precision, no need for any environmental prior, and rich applicable scenarios. However, visual-only sensor is easily affected by external environmental and motion factors, such as dynamic landmarks, the lack of texture, lighting conditions, motion blur, and few common field of view (FOV) [1]. Therefore, an inertial measurement unit (IMU), which can provide relatively accurate pose estimation information without being affected by the environment in a short period of time, can be integrated to compensate for the deficiencies of visual-only sensor [2].
However, the assumption of static landmarks in visual-related SLAM still makes it have hidden localization risk in environments with dynamic landmarks [3]. The dynamic landmarks will affect the accuracy of visual positioning, and in extreme cases, they will lead to rapid deterioration of positioning. At the same time, a clean static feature map is also of great significance for Life-Long SLAM systems. This is because that the ‘ghosting effect’ caused by the dynamic landmarks in the feature map will greatly affect the localization based on this prior map.
@@
(2) The robust and self-adaptive bundle adjustment residual for dynamic candidate landmarks: A dynamic visual residual is constructed which considers the minimum projection error prioris and is used in nonlinear optimization process to reduce the impact of dynamic candidate landmarks that deviate from the motion prior.
Compared with existing methods in the VIODE dataset [8] which contains dynamic objects within different ranges, superior robust and stable localization accuracy, and good real-time performance of our approach are demonstrated. And a clean point cloud map without ‘ghosting effect’ is obtained after the post-processing. In addition, the localization accuracy is compared in the EUROC dataset [9], and the result demonstrates stable and superior performance of our approach in static environments.
@@ II-C Methods based on inertial sensor
Cui et al. [22] detected and removed dynamic landmarks according to the epipolar constraint provided by the prior of IMU pre-integration, and conducted states estimation by utilizing more accurate static landmarks. The DynaVINS [23] proposed a robust bundle adjustment system based on the IMU pre-integration regularization factor and the momentum factor to reject dynamic landmarks. It needs to perform optimization based on the regularization factor and the momentum factor before formal nonlinear optimization to obtain the weight corresponding to each landmark in order to reject dynamic landmarks which leads to the fact that it cannot operate in real time. At the same time, its method may not be suitable for other challenging scenarios in visual SLAM, such as changes in lighting and drastic movements of camera. The method in this paper is also based on the motion prior propagated by the IMU which performs well both in dynamic environments and static environments with other challenging scenarios, and has better real-time performance.
@@ III-B Minimum Projection Error and Preprocessing
represents the precision of the epipolar constraint, is the coordinate in the normalized camera coordinate system while is the coordinate in the pixel coordinate system, is the coordinate measurement combined the true value and the residual as Fig.3, is the internal parameter matrix of camera.
@@
is the minimum square projection error defined by us. is a factor that characterizes the credibility of the motion prior obtained through IMU propagation, which is defined to avoid the incorrect elimination of dynamic landmarks caused by large IMU propagation error (mainly considered as position error) or the degradation of the epipolar constraint due to an overly small translation amount. is the variance of the position error of IMU pre-integration between two frames of camera. Noted that for the pinhole camera model, we adopt the projection error on the pixel plane, while for the fisheye camera model, we use the projection unit length error on the normalized plane.
@@
(1) Elimination in preprocess: In the situation where both dynamic and static landmarks exist, the value of CV is always greater than . Sort all landmarks in descending order according to the value of , and start eliminating from the maximum value until the CV no longer decreases or the CV decreases to the target value . The selected dynamic landmarks are either directly excluded or triangulated to compare the cleanliness of the map. The elimination criterion and the recursive formulas for mean, variance and CV of remaining landmarks are as follows.
@@ IV ROBUST BUNDLE ADJUSTMENT
When a dynamic candidate landmark enters the sliding window based optimization process, it is continuously observed for times. The visual residual for these observations will consider the prior of minimum projection error which will also be updated during the iterative optimization process. Combined with the visual residuals of static landmarks, the measurement residuals of the IMU and the prior residual, the global residual to be minimized is (7). From the probabilistic perspective, the two-dimensions’ projection residual of dynamic candidate landmarks satisfies a Gaussian distribution with a covariance matrix of minimum value as .
, and are respectively the collection of static landmarks, dynamic candidate landmarks and inertial measurements in the sliding window, while , and are respectively the covariance matrix of static landmark’s projection residual, dynamic candidate landmark’s projection residual and IMU pre-integration [24] between the and frames. represents prior information. represents the sum of cameras’ state (position,velocity and rotation at the moment ), IMU state (biases of IMU at the moment ) and landmarks’ state (the inverse depths of all the landmarks in collection and ). , and represent the IMU observation, the observation of the static landmark, and the observation of the dynamic candidate landmark respectively. And is related to the prior of minimum projection error of the observation (the form is similar to formula 3 without considering ) which is set to be when it’s value is smaller than ) obtained through the epipolar constraint precision between the first observation and the observation of dynamic candidate landmark in the sliding window. is huber kernel function [26]. Noted that .
Fig 6 shows the comparison between the original visual projection residual and the visual projection residual proposed in this paper for dynamic candidate landmarks. After optimization, 3D-2D projection from the first observation to other relevant observations is performed to calculate the reprojection errors of dynamic candidate landmarks. Remove them outside the sliding window and mark them as confirmed dynamic landmarks if the errors are still relatively large. It should be noted that during the whole process, some unstable landmarks, such as landmarks that are easily occluded, landmarks on the leaves, will also be excluded from the sliding window due to the fact that their tracking results deviate from the motion prior. Finally, a static map which is clean and free of ’ghosting effect’ will be obtained.
@@ V EXPERIMENTAL RESULTS
To evaluate our proposed algorithm, comparisons with the SOTA algorithms in the VIODE and EUROC datasets are carried out, such as VINS-FUSION [14], VINS-Ransac, VINS-Seg and Dyna-VINS [23] in the stereo mode. VINS-Seg is only tested in the VIODE dataset, since it provides instance segmentation templates which can be used to eliminate the landmarks on the dynamic vehicle instances.
@@ V-B Remove dynamic landmarks
After optimization based on sliding window using our visual residual for dynamic candidate landmarks and post-processing to them, confirmed dynamic landmarks will be removed from sliding window while others will be marked as static landmarks. The final labels of landmarks in Fig.4 are displayed in Fig.8. And the clean map of static landmarks are displayed in Fig.1.
V-C Quantitative Analysis of Localization
@@
This subsection provides a quantitative analysis of localization for our method, evaluating the localization accuracy of camera poses and the time taken. Tables I and II present the comparison of our proposed method with other state-of-the-art SLAM methods, using the RMSE (Root Mean Square Error) of ATE (Absolute Trajectory Error) and average tracking and optimization time cost as the evaluation metric. To comprehensively validate the fitness for environments within different dynamic ranges and the effectiveness of the proposed method, we conducted tests on the VIODE dynamic dataset and the EUROC static dataset.
@@
Table I Comparison of Localization accuracy with state-of-the-art methods (RMSE of ATE in [unit less]). Noted that x represents failure case (diverged).
Method
VIODE city
VIODE parking_lot
EUROC
none
low
mid
high
none
low
mid
high
V1_03
V2_03
MH_03
VINS-Fusion
0.158
0.169
0.193
0.211
0.103
0.126
0.149
0.166
0.216
0.352
0.290
VINS-Ransac
0.318
0.289
0.298
0.301
0.099
0.097
x
0.310
0.109
0.370
0.266
VINS-Seg
0.158
0.152
0.131
0.247
0.108
0.108
0.125
0.140
Dyna-VINS
0.157
0.140
0.149
0.122
0.115
0.099
0.126
0.124
0.175
0.479
0.349
IDY-VINS
0.148
0.145
0.135
0.126
0.092
0.089
0.111
0.088
0.143
0.229
0.276
@@
The effects of the proposed method on the localization accuracy is analyzed as shown in Table I. And Table II shows the analysis of our method on the tracking and optimization time cost. The VINS-Seg method does not participate in time test because it’s unable to estimate the time required for instance segmentation.
@@
(1) The improvement of localization accuracy for different environments: According to Table I, our method has a better result than VINS-Fusion and Dyna-VINS. The RANSAC method has the problem of erroneously removing static inliers when facing large-scale dynamic scenes, like Fig.4(a), which leads to the decline and instability of the accuracy. The method (VINS-Seg) which considers dynamic semantics has not achieved the most ideal localization accuracy results. Especially in the day_high sequence which contains many scenes with large-area occlusions by dynamic vehicles, it has led to a deterioration in localization. The fact indicates that directly abandon the constraints of all dynamic landmarks may be too arbitrary. Some dynamic landmarks with slow movement can still provide certainly effective constraints for the system. For city sequences of VIODE from day_none to day_high, our method improves the localization accuracy by 6.3%, 14.2%, 30.1% and 40.3% compared to VINS-Fusion. And our method improves the localization accuracy by 10.7%, 29.4%, 25.5% and 47.0% compared to VINS-Fusion for indoor parking lot sequences of VIODE from parking_lot_none to parking_lot_high. The results of EUROC sequences in real world prove that our method still has a stable improvement effect on the localization accuracy in static environments. Moreover, Dyna-VINS method may not be suitable for handling situations such as drastic camera movements and lighting changes in static environments.
@@
IDY-VINS | 19.936+0.467 | 20.397+0.362 | 20.057+0.571 | 19.982+0.926 | 38.375 | 37.375 | 37.251 | 35.564 |
(2) Better time cost performance: According to Table II, our method does not cause too much additional tracking time cost compared to RANSAC method. However, it should be noted that the time consumption will surely be higher than VINS-Fusion and Dyna-VINS which do not conduct additional elimination operations during the feature tracking phase. On the other hand, the optimized results may converge more quickly for our method than VINS-Fusion and VINS-Fusion_ransac methods because of the elimination of obvious dynamic landmarks and the classification of dynamic candidate landmarks and static landmarks during optimization process. And the more dynamic objects there are in the environment, the better improvement on the consumption of optimization time will be for our method, while the RANSAC method often requires a longer optimization time to achieve convergence with a lower localization accuracy because of incorrect elimination. Meanwhile, Dyna-VINS method requires much more time for optimization due to its complex processing. The above two aspects of tracking time cost and optimization time cost demonstrate that our method generally has a better real-time performance.
V-D Qualitative Analysis of Mapping
This subsection provides a qualitative analysis of mapping for our method. Firstly, this paper verifies the recognition and elimination effects of our algorithm on dynamic landmarks under different densities. Taking the day_high sequence as an example, Fig.9 shows the effects of the elimination of dynamic landmarks under low, medium and high mapping densities represented by the landmark densities of 120, 300 and 600 landmarks per frame respectively. The results show that our algorithm has good robustness to different mapping densities. With an increase in mapping density, the recognition and elimination effects of dynamic landmarks and unstable outliers are not attenuated, and relatively clean static mapping results can be obtained.
Secondly, with the same mapping density (600 landmarks per frame for the city sequences and 300 landmarks per frame for the parking_lot sequences of VIODE), this paper conducted mapping tests on different levels of dynamics from low to high. The results are shown in Fig.10, which verify that our algorithm can effectively recognize dynamic landmarks and leave clean static mapping results under different levels of dynamics, compared to static mapping results of VINS-Seg which abandons all landmarks on dynamic vehicles. The results of VINS-Seg in day_high and parking_lot_high sequences are shown in Fig.11.
@@ VI CONCLUSIONS
In this paper, the robust visual-inertial motion prior SLAM system (IDY-VINS) has been proposed, which effectively handles dynamic landmarks using inertial motion prior for dynamic environments to varying degrees. It is a robust system that preprocesses potential dynamic landmarks and constructs a robust and self-adaptive bundle adjustment residual for dynamic candidate landmarks to reduce their impact. After post-processing, a clean static map is obtained. Experimental evidences demonstrate that our algorithm has a superior performance compared to other algorithms in both simulated and real environments, both dynamic and static environments.
In future work, we plan to more reasonably arrange the constraints of dynamic landmarks with fully consideration of semantic information and IMU motion prior to improve the system’s localization accuracy and mapping effect. Furthermore, we plan to extend the consideration of dynamic landmarks and candidate landmarks to the overall relocalization and loop closure detection processes of VI-SLAM, so that it can provide a feasible assistance for Life-Long VI-SLAM.